OpenStudy (anonymous):

FV= PMT (1+i)n-1 **n is an exponent ________ i FV= 396 (1+0.005)72-1 _________ 0.005 FV=396 (1.005)72-1 _________ 0.005 FV= 34218 (rounded to nearest dollar) GOING BACK TO RULES OF OPERATION (PEMDAS) how was this computed, i'm still getting it wrong

OpenStudy (anonymous):

How did you get 34218?

OpenStudy (tkhunny):

1.005^72 = 1.4320442784916543292912103006685 That less one is 0.4320442784916543292912103006685 That divided by 0.005 is 86.408855698330865858242060133696 That multiplied by 396 is 34217.906856539022879863855812944 Maybe it's not your arithmetic. What's the original problem statement?

OpenStudy (anonymous):

so FV = 396(((1+.005)^72-1))/.005) Is this what you are doing?

OpenStudy (anonymous):

that was the answer the computer gave after my incorrect answer. a couple deposits $396 per month into an account paying 6% compounded monthly. Payments are made at the end of each period, how much will be in the account in 6 years? so I know n= 12*6 =72 \[\frac{ 6% }{ 12*100%? }\] (6% and 100%) it forgot the % sign + 0.005 so FV= $396 (1+0.005)72-1 ________________ WITH 72 being an exponent 0.005

OpenStudy (anonymous):

OK so first: (1+.005) first thing you do it equals 1.005 Then you do 1.005^72-1 = .4320442785 Then you do .4320442785/.005 = 86.4088557 Then you do 86.4088557 * 396 = 34217.90686 Which is FV = $34,218

OpenStudy (anonymous):

oh wow.. thank you for breaking that down.. i've been stumped for like 10 minutes

OpenStudy (anonymous):

No worries, wait till Calculus. lol. Good Luck.